40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1803     Entry time: Wed Jul 29 11:58:57 2009
Author: Alberto 
Type: Update 
Category: PSL 
Subject: MC not locking 

This morning I found the Mode Cleaner unlocked.

I check the sliders for the mirrors bias and they have not changed. Also the OSEMs readbacks show no change in the optics positions.

I don't understand what's wrong because in the previous days, in this state of alignmanet, it could lock.

I tried to tweak a little bit the periscope to check whether it was a problem of beam matching but that didn't help the cavity to lock.

I don't want to change the periscpe alignment to much becasue I believe it is still good and I suspect that there is something else going on.

ELOG V3.1.3-